OpenStudy (anonymous):

[will give medal]What is the median of the data? 19, 18, 12, 22, 15, 14, 16, 16, 14, 14, 21, 18 A. 14 B. 16 C. 17 D. 15

OpenStudy (anonymous):

put the numbers in order 12,14,14,14,15,16,16,18,18,19,21,22 Now find the middle number (the median). However, if there is an even number of numbers, there will be 2 middle numbers. So you just add the 2 middle numbers and divide by 2 to find the median. add the 2 middle numbers....then divide by 2 ...(16 + 16)/2 = 32/2 = 16 (answer B)